The radius of curvature of the concave mirror is 60 cm. At what distance from the mirror should the object be placed so that the rays that came from the object to the mirror are reflected from the mirror, and then travel parallel to each other?(answer:30 cm)


1
Expert's answer
2021-04-22T09:30:15-0400

Solution :-


Given

ROC = 60 cm


object be placed so that the rays that came from the object to the mirror are reflected from the mirror, and then travel parallel to each other

= 30 cm at focus


as we can see in ray diagram and also by mirror formula


mirror formula = "\boxed{\frac{1}{v}+\frac{1}{u}=\frac{1}{f}}"


when u = f

then V = "\infin"

"\implies"

rays will be parallel to incident rays


it will be more clear by diagram

as given below